Assume we fit an ARMA(2,0) model to the daily return on market portfolio. Model outcome is provided below.

Rt = 0.75+0.6*Rt-1+0.3* Rt-2

Given that today is Monday and return on market portfolio is 1% today and it was 0.5% yesterday, what is your forecast for Friday?
Answer #1

Given,

Rt = 0.75 + 0.6*Rt-1 + 0.3* Rt-2

RMonday = 1

RSunday = 0.5

=> RTuesday = 0.75 + 0.6*RMonday + 0.3* RSunday = 0.75 + 0.6*1+ 0.3*0.5 = 1.5

RWednesday = 0.75 + 0.6*RTuesday + 0.3* RMonday = 0.75 + 0.6*1.5+ 0.3*1 = 1.95

RThursday = 0.75 + 0.6*RWednesday + 0.3* RTuesday = 0.75 + 0.6*1.95+ 0.3*1.5 = 2.37

RFriday = 0.75 + 0.6*RThursday + 0.3* RWednesday = 0.75 + 0.6*2.37+ 0.3*1.95 = 2.757%